How many 3 letter and 3 number licenses plates can be made if you cant use the numbers 1,2,3 and you cant use the letters t,v,w,x,y and no number or letter can repeat in the plate.
Answer by stanbon(48530)

How many 3 letter and 3 number licenses plates can be made if you cant use the numbers 1,2,3 and you cant use the letters t,v,w,x,y and no number or letter can repeat in the plate.
------------------
There are 7 digits you can use.
# of ways to pick 3 of the 7 digits: 7C3
-----------------------------------
There are 21 letters you can use.
number of ways you can pick 3 of the 21 letters: 21C3
------------------------------------
# of ways to arrange the 7 objects you pick: 7!
-----------------------------------
Total # of license plates: 7C3 * 21C3 * 7! = 35*1330*5040 = 234,612,000